OpenStudy (anonymous):

I need help with my CC Algebra Sequence HW. Find the next 4 four terms in the sequence -1,-2,-3,-4..... Of course that is -5,-6,-7,-8 right? So how do I write the explicit formula Find a62 find f(10) Please help me I am stressing out so much over my homework

OpenStudy (amistre64):

is it arithmetic, geometric, or other?

OpenStudy (anonymous):

arithmetic

OpenStudy (amistre64):

good, we can either use the formula for it, or develop the formula for it

OpenStudy (amistre64):

we can see that:\[f(n)=f(n-1)-1\]it is given to us that \[f(0)=given\] \[f(1)=f(0)-1\] \[f(2)=f(1)-1\\~~~~~~=f(0)-1-1\\ ~~~~~~=f(0)-2\] \[f(3)=f(2)-1\\~~~~~~=f(0)-2-1\\ ~~~~~~=f(0)-3\] the pattern i see developing is explicitly defined as: \[f(n)=f(0)-n\]

OpenStudy (amistre64):

well, f(0) = -1, so the explicit should use that instead of f(0)

OpenStudy (amistre64):

but then different textbooks start n at different points. i started at n=0, many texts like to start n at 1, which is just a shift

OpenStudy (amistre64):

f(n) = -1 - (n-1) is one way many texts would define it

OpenStudy (anonymous):

So how would I solve the problem, Its very difficult for me to understand. I also very much appreciate you helping me

OpenStudy (amistre64):

first, figure out where you are expected to start n from; 0 or 1 then its just a matter of working thru develop a formula that works. if n starts at 1; it is easy to see that f(1) = -1 f(2) = -2 f(3) = -3 ... f(n) = -n if n starts at 0; it is easy as well to see that f(0) = -1 f(1) = -2 f(2) = -3 f(3) = -4 ... f(n) = -(n+1)

OpenStudy (anonymous):

Again Its very hard for me to understand it, Is there any way you can put it to help me better.

OpenStudy (anonymous):

In class I knew how to do it, and even in extra help I did. now for some reason I seemed to forget

OpenStudy (anonymous):

I just need like a refresher if you know what I mean

OpenStudy (amistre64):

i dont think so, it is rather simplistic to me and I cannot see where you would be having issues at. Can you tell me what is confusing you?

OpenStudy (anonymous):

How do I find the the explicit formula for my problem

OpenStudy (amistre64):

find a rule or pattern that is being expressed by the sequence. For this particular sequence of numbers that pattern is hopefully self explanatory. If it is form that is confusing: implicit defines each new term by the other terms, but has to define some initial terms to start with:\[f(n)=f(n-1)-2f(n-2)\\f(0)=1,~f(1)=7\]using this information we can then find f(2,3,4,5,...) an explicit formula is a form that simply find the nth term by using the value of n \[f(n)=2^n+3n\]this way we do not need to know any initial values since any value of n defines the nth term

OpenStudy (amistre64):

the pattern that shows up in your problem is easy to define explicitly: f(n) = -n ; assuming n starts at 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Let me try it, and ill let you know if I need any more assistance

OpenStudy (anonymous):

Thanks you so much !!

OpenStudy (amistre64):

i might have to go home soon so if im not around, feel free to ask others as well :)

OpenStudy (anonymous):

Ok thank I will let you know